# Medium effects on the electromagnetic form factors of $\rho$ meson

### v1

@@ V In-medium $\rho$ meson EMFFs
Here we describe the $\rho$ meson EMFFs in the nuclear medium in the covariant NJL model. Shortly, the generic expression for the electromagnetic current for the $\rho$ meson can be parameterized in terms of three form factors,
@@
where the subscript $i=\big{(}\rho,\omega\big{)}$ is the dressed quark form factor obtained from the inhomogeneous BSE. Note that $F^{*}_{1\omega}(Q^{2})=F^{*}_{1\omega}(Q^{2})=$ 1 and $F^{*}_{2\rho}(Q^{2})=F^{*}_{2\omega}(Q^{2})=$ 0, for the pointlike quark. This implies $F_{2Q}^{*}(Q^{2})=\frac{1}{6}F_{2\omega}^{*}(Q^{2})\pm\frac{1}{2}F_{2\rho}^{*}(Q^{2})=0$ where $Q=(U,D)$ (dressed quarks). However, $f_{1}^{T*}(Q^{2}),f_{2}^{T*}(Q^{2})$, and $f_{3}^{T*}(Q^{2})$ are not zero as can be seen in Refs. [13, 54]. It is worth noting that $f_{3}^{T}(Q^{2})=0$ as it preserves charge conservation.
@@ VI Numerical Result
Here, we present the numerical results for the $\rho$ meson properties and their EMFFs, such as the charge radius $\big{<}r_{\rho}^{*}\big{>}\equiv\big{<}r^{*2}_{C}\big{>}^{1/2}$ (See Eq. (47)), the charge form factor $G_{C}^{*}(Q^{2})$, magnetic form factor $G_{M}^{*}(Q^{2})$, and quadrupole moment form factor $G_{Q}^{*}(Q^{2})$) in the nuclear medium. The NJL model parameter used in the present work is $G_{\pi}$, $G_{\omega}$, $G_{\rho}$, $G_{a}$, and $G_{s}$, as used in Refs. [34, 33, 53]. In this work, the regularization parameter $\Lambda_{\mathrm{IR}}$ in the Schwinger proper-time regularization scheme is fixed at 240 MeV, which is comparable to $\Lambda_{\mathrm{QCD}}$, while the free space (vacuum) constituent quark mass is taken as $M_{0}=400$ MeV for $u$ and $d$ quarks.
The remaining free space parameters are fitted to the physical pion mass $m_{\pi}=$ 140 MeV, physical kaon mass $m_{K}=$ 495 MeV, nucleon mass $M_{N}=M_{N0}=$ 940 MeV, and $\rho$ meson mass $m_{\rho}=$ 770 MeV, together with the pion weak-decay constant $f_{\pi}=$ 93 MeV. The fitted results give an ultraviolet cutoff $\Lambda_{\mathrm{UV}}=$ 645 MeV, pion coupling constant $G_{\pi}=$ 19.04 $\times$ $10^{-6}$ MeV-2, axial-vector diquark coupling constant $G_{a}=$ 2.8 $\times$ $10^{-6}$ MeV-2, scalar diquark coupling constant $G_{s}=$ 7.49 $\times$ $10^{-6}$ MeV-2, and strange constituent quark mass $M_{s}=$ 611 MeV. The $\phi$ meson mass is taken $m_{\phi}=$ 1001 MeV, scalar diquark mass $M_{s}=$ 687 MeV, and axial diquark mass $M_{a}=$ 1027 MeV. It is worth noting that the coupling constants $G_{s}$ and $G_{a}$ are determined by solving the Faddeev equation to reproduce the nucleon mass in vacuum, $M_{N}$, and axial coupling constant of nucleon $g_{A}=1.267$, as employed in Ref. [53].
The parameters for symmetric nuclear matter are determined by fitting the (negative of) energy per nucleon, $E_{B}/A=-15.7$ MeV, at the saturation baryon density $\rho_{0}=0.16$ fm-3. This yields a value of $G_{\omega}=6.03\times 10^{-6}$ MeV-2. The current quark mass value (in vacuum) for the up quark (and down quark) is taken as $m_{u}=m_{d}=16$ MeV (SU(2) isospin symmetry), while the current mass of the strange quark is set to $m_{s}=356$ MeV, following Refs. [34, 33, 53].
@@ VI.1 In-medium $\rho$ meson properties
Results for the $\rho$ meson properties in nuclear medium are shown in Fig. 2(a)-(d). Evaluating the pole position of the $t$-matrix in Eq. (22), we obtain the $\rho$ meson mass in nuclear medium, as depicted in Fig. 2(a). It shows that the $\rho$ meson mass decreases as the nuclear matter density increases. A similar behavior was found in Refs. [52, 58], represented by the filled asterisk (red) data, where the in-medium $\rho$ meson mass also decreases as the nuclear matter density increases. However, the magnitude is significantly different, in particular at $\rho_{B}/\rho_{0}\gtrsim$ 0.5. It is worth noting that in Ref. [52], they used $\rho_{0}$ is 0.15 fm-3, which is different from the present work, $\rho_{0}=0.16$ fm-3. The difference in the $\rho$ meson mass in the nuclear medium is more pronounced at higher densities due to the strong attractive scalar potential in the QMC model, as clearly indicated in Fig. 2(a). We find that the $\rho$ mass reduction relative to that in free space is approximately 10% at normal nuclear matter density $\rho_{B}/\rho_{0}=1.00$, which also may correspond to the amount of the chiral symmetry restoration. Our result is consistent with that obtained by the QCDSR of Ref. [43] and a very recent analysis in the IQCDSR of Ref. [45], which are about (10-20)% in the $\rho$ mass reduction. However, our result in $\rho$ mass reduction is rather different from that obtained in Ref. [52], which is approximated by 33%. The difference is probably related to a larger constituent light quark mass $m_{q}$ used in the model of Ref. [52] (See Table I of the reference), where this value of $m_{q}=430$ MeV was set to be able to have the ”bound $\rho$ meson state”, yielding the larger $g_{\sigma}$ and $g_{\omega}$ coupling constant values to reproduce the empirically extracted nuclear matter saturation properties in the QMC model.
@@
Figure 2: Results for (a) $\rho$ meson effective mass, (b) dynamical quark effective mass, (c) $\rho$-quark coupling constant, and (d) nucleon effective mass as a function of $\rho_{B}/\rho_{0}$, where $\rho_{B}$ and $\rho_{0}\,\,(=0.16\,\,\text{fm}^{-3})$ are respectively the baryon and saturation densities. Note that the in-medium $\rho$ meson and dynamical effective quark masses represented by the filled asterisk (red) data are taken from Ref. [52]. Note that, as mentioned in the text, the results of Ref. [52] used $\rho_{0}=0.15$ fm-3 and thus, for an exact comparison with the present results ($\rho_{0}=0.16$ fm-3), we adjust the horizontal values of the filled asterisk (red) data multiplying by a factor of (0.15/0.16) = 0.9375. Additionally, the ”hybrid” model of the $\rho$-meson in Ref. [52] can form the bound state $\rho$ meson up to around $\rho_{B}/\rho_{0}\simeq 0.9\times 0.9375=0.84375$ in the present $\rho_{B}/\rho_{0}$ units. Thus, for larger nuclear matter densities, the in-medium $\rho$ meson properties could not be calculated within the ”hybrid model” in Ref. [52].
@@
In Fig. 2(b), we also show the dynamical quark effective mass in symmetric nuclear matter calculated in the NJL model that was described in Sec. III. It is worth noting that this in-medium dynamical quark mass was used as input to calculate the $\rho$ meson mass in symmetric nuclear matter within the NJL model. As in the features of the in-medium $\rho$ meson mass, the dynamical quark mass also decreases with increasing nuclear matter density. The trend of our results for the $\rho$ meson and dynamical quark masses is consistent with other theoretical predictions [52], but different in magnitude, as clearly indicated in Fig. 2(b). We also compare our present result with that calculated in the QMC model reported in Ref. [52]. We find that the differences are more significant at higher densities. Again, these differences are expected due to a strong scalar attractive potential originated from the QMC model. We find that the reduction of the dynamical quark effective mass at $\rho_{B}/\rho_{0}=1.0$ relative to that in free space (vacuum) is approximately 18% (72 MeV). The reduction of the constituent quark mass at $\rho_{B}/\rho_{0}=0.9$ obtained in Ref. [52] is 40% (170 MeV), which is larger than our result. Note that to be exactly comparable with the present unit scale of $\rho_{0}=$ 0.16 fm-3, the nuclear matter density has to be adjusted by a factor of $0.9375$, for example, $\rho_{B}/\rho_{0}=0.9\times(0.15/0.16)=0.9\times 0.9375=0.84375$, as shown in Figs. 2(a)-(b).
@@
Figure 2(d) shows the results of the nucleon effective mass as a function of the nuclear matter density. The nucleon effective mass is calculated in the NJL nuclear matter model described in Sec. III, which shows the unique features of the NJL model approach computed at the quark level. As a result, it shows that, as the nuclear matter density increases, the nucleon effective mass decreases, which is consistent with other theoretical predictions [52, 59]. The reduction of the nucleon effective mass at normal nuclear matter density relative to that in free space is about 17% (156 MeV). We note that the NJL model approach to nuclear matter reproduces the (negative of) energy for nuclear matter at saturation density $\rho_{0}=$ 0.16 fm-3, as explained in Ref. [53]. This indicates that the NJL model description of nuclear matter guarantees the nuclear matter stability, as empirically proven.
@@ VI.2 In-medium $\rho$ meson EMFFs
Before presenting the results for the charge, magnetic moment, and quadrupole form factors of the $\rho^{+}$ meson in symmetric nuclear matter as well as in free space, we explore first the results of the BSE dressed up-quark form factors [^2] In Ref. [13], $F_{\mathrm{1U}}^{*}(Q^{2})$ and $F_{\mathrm{1D}}^{*}(Q^{2})$ are usually written as $F_{\mathrm{1U}}^{*\mathrm{BSE}}(Q^{2})$ and $F_{\mathrm{1D}}^{*\mathrm{BSE}}(Q^{2})$, respectively, where the BSE superscript indicates that the form factors are uniquely determined in the BSE. In this work, we suppress it for simplicity. $F_{\mathrm{1U}}^{*}(Q^{2})$ and down-quark (antidown-quark can be obtained by multiplying by a factor of 1-1) form factors, $F_{\mathrm{1D}}^{*}(Q^{2})$, where the expression of this inhomogeneous BSE dressed quark form factors are formulated in Eq. (41).
@@
Results for BSE dressed down-quark form factor $F_{\mathrm{1D}}^{*}(Q^{2})$ for different nuclear matter densities as a function of $Q^{2}$ are shown in Fig. 3(a), where we can see that $F_{\mathrm{1D}}^{*}(Q^{2})$ increases as the nuclear matter density increases. At $Q^{2}=0$ we have $F_{\mathrm{1D}}^{*}(Q^{2}=0)=e_{d}=-1/3$, where $e_{d}$ stands for the electric charge of the down quark. It is worth noting that this normalization value at $Q^{2}=0$ is normalized not only in free space but also in nuclear matter. Also, we find that $F_{\mathrm{1D}}^{*}(Q^{2})$ increases up to around $Q^{2}\simeq$ 2 GeV2 and begins to decrease as the $Q^{2}$ increases. It is worth noting that, in this work, we do not consider the pion loops contributions in the $\rho$ meson EMFFs as in Refs. [13, 54], since we focus on the nuclear medium effects on the $\rho$ meson EMFFs. The pion loop contribution to the $\rho$ meson EMFFs in nuclear matter is expected to be insignificant, because the physical pion mass remains nearly unchanged in the nuclear medium, as indicated in Ref. [31] (and references therein).
As $Q^{2}\rightarrow\infty$ in free space, the $F_{\mathrm{1D}}^{*}(Q^{2})$ does not go to zero, but it behaves as $F_{\mathrm{1D}}^{*}(Q^{2})\simeq e_{d}$. A similar feature is captured in the $F_{\mathrm{1D}}^{*}(Q^{2})$ in nuclear matter, where it also gives $F_{\mathrm{1D}}^{*}(Q^{2})\simeq e_{d}$ at $Q^{2}\rightarrow\infty$. Such a behavior at larger $Q^{2}$ is consistent with the expectation results from asymptotic freedom of QCD [54]. We also compute the BSE dressed form factors for the up quark. The results for the $F_{\mathrm{1U}}^{*}(Q^{2})$ as a function of $Q^{2}$ for different nuclear matter densities are illustrated in Fig. 3(b). It is clearly shown that the behavior of $F_{\mathrm{1U}}^{*}(Q^{2})$ is rather different from that for $F_{\mathrm{1D}}^{*}(Q^{2})$, where the $F_{\mathrm{1U}}^{*}(Q^{2})$ decreases as the nuclear matter density increases. As a function of $Q^{2}$, the $F_{\mathrm{1U}}^{*}(Q^{2})$ form factor decreases up to $Q^{2}\simeq 2$ GeV2, and then increases at higher $Q^{2}$, exhibiting an opposite trend to that of $F_{\mathrm{1D}}^{*}(Q^{2})$. It is worth noting that if the absolute value is taken, or the sign is changed for the dressed $D$ quark form factor as $F_{\mathrm{1D}}^{*}(Q^{2})$, the global behavior looks similar except for the magnitude difference with $F_{\mathrm{1U}}^{*}(Q^{2})$.
@@
Figure 4(a) shows that the $f_{1}^{*V}(Q^{2})$ decreases as the nuclear matter density $\rho_{B}/\rho_{0}$ and $Q^{2}$ increase. Note that the $f_{1}^{V}(Q^{2})$ is unity at $Q^{2}=0$ in free space as well as in nuclear matter. This corresponds that $f_{1}^{*V}(Q^{2})$ satisfies the charge conservation. Such a normalization condition is also found in Refs. [13, 54] for a free space case. Furthermore, Fig. 4(a) also exhibits that the change of $f_{1}^{*V}(Q^{2})$ is more pronounced at higher densities.
In addition to the results for $f_{1}^{*V}(Q^{2})$, we also show the results for the vector body form factors $f_{2}^{*V}(Q^{2})$ for different nuclear matter densities as a function of $Q^{2}$, as illustrated in Fig. 4(b). We find that the $f_{2}^{*V}(Q^{2})$ also decreases as the $\rho_{B}/\rho_{0}$ and $Q^{2}$ increase. Also, the values of $f_{2}^{*V}(Q^{2}=0)$ decrease as the nuclear matter density increases. The reduction of $f_{2}^{*V}(Q^{2}=0)$ at $\rho_{B}/\rho_{0}=1.0$ is 6.87% in comparison to that for the free space. At higher density $\rho_{B}/\rho_{0}=2.0$, the decrease of $f_{2}^{*V}(Q^{2}=0)$ is 10.49%. In Fig. 4(c), we show our results for $f_{3}^{*V}(Q^{2})$ as a function of $Q^{2}$ for different densities. In free space, $f_{3}^{*V}(Q^{2}=0)=$ 2.09, which is the magnetic moment of the $\rho^{+}$ meson in the units of $e/(2m_{\rho})$, where $e=1$ and $m_{\rho}$ are the natural units of the positron charge $e$ and $\rho$ meson mass, respectively. This value is rather larger than the canonical value [54]. In symmetric nuclear matter, the magnetic moment decreases as the nuclear matter density increases. For example, the values of $f_{3}^{*V}(Q^{2}=0)=2.079$ at $\rho_{B}/\rho_{0}=1.0$. For $\rho_{B}/\rho_{0}=2.0$, we have $f_{3}^{*V}(Q^{2}=0)=2.077$.

### v2

@@ V In-medium $\rho$ meson EMFFs
Here we describe the $\rho$ meson EMFFs in the nuclear medium in the covariant NJL model. Shortly, the generic expression for the electromagnetic current for the $\rho$ meson can be parametrized in terms of three form factors,
@@
where the subscript $i=\big(\rho,\omega\big)$ is the dressed quark form factor obtained from the inhomogeneous BSE. Note that $F^{*}_{1\omega}(Q^{2})=F^{*}_{1\omega}(Q^{2})=$ 1 and $F^{*}_{2\rho}(Q^{2})=F^{*}_{2\omega}(Q^{2})=$ 0, for the pointlike quark. This implies $F_{2Q}^{*}(Q^{2})=\frac{1}{6}F_{2\omega}^{*}(Q^{2})\pm\frac{1}{2}F_{2\rho}^{*}(Q^{2})=0$ where $Q=(U,D)$ (dressed quarks). However, $f_{1}^{T*}(Q^{2}),f_{2}^{T*}(Q^{2})$, and $f_{3}^{T*}(Q^{2})$ are not zero as can be seen in Refs. [Carrillo-Serrano:2015uca, Cloet:2014rja]. It is worth noting that $f_{3}^{T}(Q^{2})=0$ as it preserves charge conservation.
@@ VI Numerical Result
Here, we present the numerical results for the $\rho$ meson properties and their EMFFs, such as the charge radius $\big<r_{\rho}^{*}\big>\equiv\big<r^{*2}_{C}\big>^{1/2}$ (See Eq. (47)), the charge form factor $G_{C}^{*}(Q^{2})$, magnetic form factor $G_{M}^{*}(Q^{2})$, and quadrupole moment form factor $G_{Q}^{*}(Q^{2})$) in the nuclear medium. The NJL model parameter used in the present work is $G_{\pi}$, $G_{\omega}$, $G_{\rho}$, $G_{a}$, and $G_{s}$, as used in Refs. [Hutauruk:2021kej, Hutauruk:2016sug, Bentz:2001vc]. In this work, the regularization parameter $\Lambda_{\mathrm{IR}}$ in the Schwinger proper-time regularization scheme is fixed at 240 MeV, which is comparable to $\Lambda_{\mathrm{QCD}}$, while the free space (vacuum) constituent quark mass is taken as $M_{0}=400$ MeV for $u$ and $d$ quarks.It is worth mentioning that the fixing of the dynamical quark mass in the NJL model is somewhat arbitrary; however, the results obtained for other value choices of $M_{0}$ are qualitatively expected to be quite similar, as indicated in Ref. [Bentz:2001vc].
The remaining free space parameters are fitted to the physical pion mass $m_{\pi}=$ 140 MeV, physical kaon mass $m_{K}=$ 495 MeV, nucleon mass $M_{N}=M_{N0}=$ 940 MeV, and $\rho$ meson mass $m_{\rho}=$ 770 MeV, together with the pion weak-decay constant $f_{\pi}=$ 93 MeV. The fitted results give an ultraviolet cutoff $\Lambda_{\mathrm{UV}}=$ 645 MeV, pion coupling constant $G_{\pi}=$ 19.04 $\times$ $10^{-6}$ MeV-2, axial-vector diquark coupling constant $G_{a}=$ 2.8 $\times$ $10^{-6}$ MeV-2, scalar diquark coupling constant $G_{s}=$ 7.49 $\times$ $10^{-6}$ MeV-2, and strange constituent quark mass $M_{s}=$ 611 MeV. The $\phi$ meson mass is taken $m_{\phi}=$ 1001 MeV, scalar diquark mass $M_{sdi}=$ 687 MeV, and axial diquark mass $M_{adi}=$ 1027 MeV. It is worth noting that the coupling constants $G_{s}$ and $G_{a}$ are determined by solving the Faddeev equation to reproduce the nucleon mass in vacuum, $M_{N}$, and axial coupling constant of nucleon $g_{A}=1.267$, as employed in Ref. [Bentz:2001vc].
The parameters for symmetric nuclear matter are determined by fitting the (negative of) energy per nucleon, $E_{B}/A=-15.7$ MeV, at the saturation baryon density $\rho_{0}=0.16$ fm-3. This yields a value of $G_{\omega}=6.03\times 10^{-6}$ MeV-2. The current quark mass value (in vacuum) for the up quark (and down quark) is taken as $m_{u}=m_{d}=16$ MeV (SU(2) isospin symmetry), while the current mass of the strange quark is set to $m_{s}=356$ MeV, following Refs. [Hutauruk:2021kej, Hutauruk:2016sug, Bentz:2001vc].
@@ VI.1 In-medium $\rho$ meson properties
Results for the $\rho$ meson properties in nuclear medium are shown in Fig. 2(a)-(d). Evaluating the pole position of the $t$-matrix in Eq. (22), we obtain the $\rho$ meson mass in nuclear medium, as depicted in Fig. 2(a). It shows that the $\rho$ meson mass decreases as the nuclear matter density increases. A similar behavior was found in Refs. [deMelo:2018hfw, Shakin:1993ta], represented by the filled asterisk (red) data, where the in-medium $\rho$ meson mass also decreases as the nuclear matter density increases. However, the magnitude is significantly different, in particular at $\rho_{B}/\rho_{0}\gtrsim$ 0.5. It is worth noting that in Ref. [deMelo:2018hfw], they used $\rho_{0}$ is 0.15 fm-3, which is different from the present work, $\rho_{0}=0.16$ fm-3. The difference in the $\rho$ meson mass in the nuclear medium is more pronounced at higher densities due to the strong attractive scalar potential in the QMC model, as clearly indicated in Fig. 2(a). We find that the $\rho$ mass reduction relative to that in free space is approximately 10% at normal nuclear matter density $\rho_{B}/\rho_{0}=1.00$, which also may correspond to the amount of the chiral symmetry restoration. Our result is consistent with that obtained by the QCDSR of Ref. [Hatsuda:1991ez] and a very recent analysis in the IQCDSR of Ref. [Mutuk:2025lak], which are about (10–20)% in the $\rho$ mass reduction. However, our result in $\rho$ mass reduction is rather different from that obtained in Ref. [deMelo:2018hfw], which is approximated by 33%. The difference is probably related to a larger constituent light quark mass $m_{q}$ used in the model of Ref. [deMelo:2018hfw] (See Table I of the reference), where this value of $m_{q}=430$ MeV was set to be able to have the ”bound $\rho$ meson state”, yielding the larger $g_{\sigma}$ and $g_{\omega}$ coupling constant values to reproduce the empirically extracted nuclear matter saturation properties in the QMC model.
@@
Figure 2: Results for (a) $\rho$ meson effective mass, (b) dynamical quark effective mass, (c) $\rho$-quark coupling constant, and (d) nucleon effective mass as a function of $\rho_{B}/\rho_{0}$, where $\rho_{B}$ and $\rho_{0}\,\,(=0.16\,\,\text{fm}^{-3})$ are respectively the baryon and saturation densities. Note that the in-medium $\rho$ meson and dynamical effective quark masses represented by the filled asterisk (red) data are taken from Ref. [deMelo:2018hfw]. Note that, as mentioned in the text, the results of Ref. [deMelo:2018hfw] used $\rho_{0}=0.15$ fm-3 and thus, for an exact comparison with the present results ($\rho_{0}=0.16$ fm-3), we adjust the horizontal values of the filled asterisk (red) data multiplying by a factor of (0.15/0.16) = 0.9375. Additionally, the ”hybrid” model of the $\rho$-meson in Ref. [deMelo:2018hfw] can form the bound state $\rho$ meson up to around $\rho_{B}/\rho_{0}\simeq 0.9\times 0.9375=0.84375$ in the present $\rho_{B}/\rho_{0}$ units. Thus, for larger nuclear matter densities, the in-medium $\rho$ meson properties could not be calculated within the ”hybrid model” in Ref. [deMelo:2018hfw].
@@
In Fig. 2(b), we also show the dynamical quark effective mass in symmetric nuclear matter calculated in the NJL model that was described in Sec. III. It is worth noting that this in-medium dynamical quark mass was used as input to calculate the $\rho$ meson mass in symmetric nuclear matter within the NJL model. As in the features of the in-medium $\rho$ meson mass, the dynamical quark mass also decreases with increasing nuclear matter density. The trend of our results for the $\rho$ meson and dynamical quark masses is consistent with other theoretical predictions [deMelo:2018hfw], but different in magnitude, as clearly indicated in Fig. 2(b). We also compare our present result with that calculated in the QMC model reported in Ref. [deMelo:2018hfw]. We find that the differences are more significant at higher densities. Again, these differences are expected due to a strong scalar attractive potential originated from the QMC model. We find that the reduction of the dynamical quark effective mass at $\rho_{B}/\rho_{0}=1.0$ relative to that in free space (vacuum) is approximately 18% (72 MeV). The reduction of the constituent quark mass at $\rho_{B}/\rho_{0}=0.9$ obtained in Ref. [deMelo:2018hfw] is 40% (170 MeV), which is larger than our result. Note that to be exactly comparable with the present unit scale of $\rho_{0}=$ 0.16 fm-3, the nuclear matter density has to be adjusted by a factor of $0.9375$, for example, $\rho_{B}/\rho_{0}=0.9\times(0.15/0.16)=0.9\times 0.9375=0.84375$, as shown in Figs. 2(a)-(b).
@@
Figure 2(d) shows the results of the nucleon effective mass as a function of the nuclear matter density. The nucleon effective mass is calculated in the NJL nuclear matter model described in Sec. III, which shows the unique features of the NJL model approach computed at the quark level. As a result, it shows that, as the nuclear matter density increases, the nucleon effective mass decreases, which is consistent with other theoretical predictions [deMelo:2018hfw, Zhu:2018vwn]. The reduction of the nucleon effective mass at normal nuclear matter density relative to that in free space is about 17% (156 MeV). We note that the NJL model approach to nuclear matter reproduces the (negative of) energy for nuclear matter at saturation density $\rho_{0}=$ 0.16 fm-3, as explained in Ref. [Bentz:2001vc]. This indicates that the NJL model description of nuclear matter guarantees the nuclear matter stability, as empirically proven.
@@ VI.2 In-medium $\rho$ meson EMFFs
Before presenting the results for the charge, magnetic moment, and quadrupole form factors of the $\rho^{+}$ meson in symmetric nuclear matter as well as in free space, we explore first the results of the BSE dressed up-quark form factors [^2] In Ref. [Carrillo-Serrano:2015uca], $F_{\mathrm{1U}}^{*}(Q^{2})$ and $F_{\mathrm{1D}}^{*}(Q^{2})$ are usually written as $F_{\mathrm{1U}}^{*\mathrm{BSE}}(Q^{2})$ and $F_{\mathrm{1D}}^{*\mathrm{BSE}}(Q^{2})$, respectively, where the BSE superscript indicates that the form factors are uniquely determined in the BSE. In this work, we suppress it for simplicity. $F_{\mathrm{1U}}^{*}(Q^{2})$ and down-quark (antidown-quark can be obtained by multiplying by a factor of 1-1) form factors, $F_{\mathrm{1D}}^{*}(Q^{2})$, where the expression of this inhomogeneous BSE dressed quark form factors are formulated in Eq. (41).
@@
Results for BSE dressed down-quark form factor $F_{\mathrm{1D}}^{*}(Q^{2})$ for different nuclear matter densities as a function of $Q^{2}$ are shown in Fig. 3(a), where we can see that $F_{\mathrm{1D}}^{*}(Q^{2})$ increases as the nuclear matter density increases. At $Q^{2}=0$ we have $F_{\mathrm{1D}}^{*}(Q^{2}=0)=e_{d}=-1/3$, where $e_{d}$ stands for the electric charge of the down quark. It is worth noting that this normalization value at $Q^{2}=0$ is normalized not only in free space but also in nuclear matter. Also, we find that $F_{\mathrm{1D}}^{*}(Q^{2})$ increases up to around $Q^{2}\simeq$ 2 GeV2 and begins to decrease as the $Q^{2}$ increases. It is worth noting that, in this work, we do not consider the pion loops contributions in the $\rho$ meson EMFFs as in Refs. [Carrillo-Serrano:2015uca, Cloet:2014rja], since we focus on the nuclear medium effects on the $\rho$ meson EMFFs. The pion loop contribution to the $\rho$ meson EMFFs in nuclear matter is expected to be insignificant, because the physical pion mass remains nearly unchanged in the nuclear medium, as indicated in Ref. [Gifari:2024ssz] (and references therein).
As $Q^{2}\rightarrow\infty$ in free space, the $F_{\mathrm{1D}}^{*}(Q^{2})$ does not go to zero, but it behaves as $F_{\mathrm{1D}}^{*}(Q^{2})\simeq e_{d}$. A similar feature is captured in the $F_{\mathrm{1D}}^{*}(Q^{2})$ in nuclear matter, where it also gives $F_{\mathrm{1D}}^{*}(Q^{2})\simeq e_{d}$ at $Q^{2}\rightarrow\infty$. Such a behavior at larger $Q^{2}$ is consistent with the expectation results from asymptotic freedom of QCD [Cloet:2014rja]. We also compute the BSE dressed form factors for the up quark. The results for the $F_{\mathrm{1U}}^{*}(Q^{2})$ as a function of $Q^{2}$ for different nuclear matter densities are illustrated in Fig. 3(b). It is clearly shown that the behavior of $F_{\mathrm{1U}}^{*}(Q^{2})$ is rather different from that for $F_{\mathrm{1D}}^{*}(Q^{2})$, where the $F_{\mathrm{1U}}^{*}(Q^{2})$ decreases as the nuclear matter density increases. As a function of $Q^{2}$, the $F_{\mathrm{1U}}^{*}(Q^{2})$ form factor decreases up to $Q^{2}\simeq 2$ GeV2, and then increases at higher $Q^{2}$, exhibiting an opposite trend to that of $F_{\mathrm{1D}}^{*}(Q^{2})$. It is worth noting that if the absolute value is taken, or the sign is changed for the dressed $D$ quark form factor as $F_{\mathrm{1D}}^{*}(Q^{2})$, the global behavior looks similar except for the magnitude difference with $F_{\mathrm{1U}}^{*}(Q^{2})$.
@@
Figure 4(a) shows that the $f_{1}^{*V}(Q^{2})$ decreases as the nuclear matter density $\rho_{B}/\rho_{0}$ and $Q^{2}$ increase. Note that the $f_{1}^{V}(Q^{2})$ is unity at $Q^{2}=0$ in free space as well as in nuclear matter. This corresponds that $f_{1}^{*V}(Q^{2})$ satisfies the charge conservation. Such a normalization condition is also found in Refs. [Carrillo-Serrano:2015uca, Cloet:2014rja] for a free space case. Furthermore, Fig. 4(a) also exhibits that the change of $f_{1}^{*V}(Q^{2})$ is more pronounced at higher densities.
In addition to the results for $f_{1}^{*V}(Q^{2})$, we also show the results for the vector body form factors $f_{2}^{*V}(Q^{2})$ for different nuclear matter densities as a function of $Q^{2}$, as illustrated in Fig. 4(b). We find that the $f_{2}^{*V}(Q^{2})$ also decreases as the $\rho_{B}/\rho_{0}$ and $Q^{2}$ increase. Also, the values of $f_{2}^{*V}(Q^{2}=0)$ decrease as the nuclear matter density increases. The reduction of $f_{2}^{*V}(Q^{2}=0)$ at $\rho_{B}/\rho_{0}=1.0$ is 6.87% in comparison to that for the free space. At higher density $\rho_{B}/\rho_{0}=2.0$, the decrease of $f_{2}^{*V}(Q^{2}=0)$ is 10.49%. In Fig. 4(c), we show our results for $f_{3}^{*V}(Q^{2})$ as a function of $Q^{2}$ for different densities. In free space, $f_{3}^{*V}(Q^{2}=0)=$ 2.09, which is the magnetic moment of the $\rho^{+}$ meson in the units of $e/(2m_{\rho})$, where $e=1$ and $m_{\rho}$ are the natural units of the positron charge $e$ and $\rho$ meson mass, respectively. This value is rather larger than the canonical value [Cloet:2014rja]. In symmetric nuclear matter, the magnetic moment decreases as the nuclear matter density increases. For example, the values of $f_{3}^{*V}(Q^{2}=0)=2.079$ at $\rho_{B}/\rho_{0}=1.0$. For $\rho_{B}/\rho_{0}=2.0$, we have $f_{3}^{*V}(Q^{2}=0)=2.077$.